OpenStudy (anonymous):

Medal, Fan, Testimony!! If AB measures 155°, what is the length of the radius of the circle, to the nearest hundredth?

OpenStudy (anonymous):



OpenStudy (anonymous):

I am sorry, but I can't help you on this one. @iGreen may be able too.

OpenStudy (campbell_st):

so there must be a little more information, like circumference or area.. or arc length..?

OpenStudy (anonymous):

 sorry forgot that

OpenStudy (campbell_st):

ok... so is 14 the arc length...?

OpenStudy (anonymous):

yes for the AB

OpenStudy (campbell_st):

ok... so then the arc length of can be found using \[l = \frac{\theta}{360} \times 2pir~~~~~~~ \theta~ is ~the ~sector ~\angle.\] so you have \[14 = \frac{155}{360} \times 2 \times \pi \times r\] you need to solve for r

OpenStudy (anonymous):

Thank you so much

OpenStudy (campbell_st):

hope it makes sense

OpenStudy (anonymous):

yes it does i have one more question

OpenStudy (campbell_st):

post it as a new question and I'll look at it